### v3

@@ Abstract
We address two important statistical problems: that of estimating for mixtures of multivariate normal distributions and mixtures of $t$-distributions based of univariate projections, and that of measuring the agreement between two different random partitions. The results are based on an earlier work of the authors, where it was shown that mixtures of multivariate Gaussian or $t$-distributions can be distinguished by projecting them onto a certain predetermined finite set of lines, the number of lines depending only on the total number of distributions involved and on the ambient dimension. We also compare our proposal with robust versions of the expectation-maximization method EM. In each case, we present algorithms for effecting the task, and compare them with existing methods by carrying out some simulations.
@@ 1.1.1 RP via the Johnson–Lindenstrauss lemma
Theorem 1 (Johnson–Lindenstrauss [46]).
Let $\aleph_{n}$ be a set of $n$ points in $\mathbb{R}^{d}$, let $0<\epsilon<1$ and let $q>(8\log n)/\epsilon^{2}$. Then there exists a linear map $f:\mathbb{R}^{d}\to\mathbb{R}^{q}$ such that, for all $u,v\in\aleph_{n}$,
@@
Let $\aleph_{n}$ and $\epsilon$ be as in Theorem 1, and let $q>(4\log n)/(\epsilon^{2}/2-\epsilon^{3}/3)$. Define $f(u):=\sqrt{(d/q)}PUu$, where $U$ is a $d\times d$ unitary matrix whose columns are iid random vectors uniformly distributed on the unit sphere in $\mathbb{R}^{d}$, and $P:\mathbb{R}^{d}\to\mathbb{R}^{q}$ denotes the projection onto the first $q$ coordinates. Then
@@ 2.3 Strong sm-uniqueness sets
Theorems 5 and 6 beg the question as to whether there exist strong sm-uniqueness sets of arbitrarily large cardinality. An affirmative answer is given in the following result based on [34, Theorem 4.3], which also provides a realistic method for generating them.
@@ 3.2 Description of the algorithm
$\sum_{j=1}^{n}{\lambda_{j}}\frac{1}{(2\pi\det(\Sigma_{j}))^{1/2}}\exp\Bigl{(}-\tfrac{1}{2}(x-\mu_{j})^{T}\Sigma_{j}^{-1}(x-\mu_{j})\Bigr{)},\quad(x\in\mathbb{R}^{d}),$
where $\lambda_{1},\dots,\lambda_{n}\in[0,1]$, $\sum_{j=1}^{n}\lambda_{j}=1$, $\mu_{1},\dots,\mu_{n}\in\mathbb{R}^{d}$, and $\Sigma_{1},\dots,\Sigma_{n}$ are $d\times d$ positive-definite matrices. We denote $\Lambda:=(\lambda_{1},\dots,\lambda_{n})$, $\bm{\mu}:=(\mu_{1},\dots,\mu_{n})$, $\bm{\Sigma}:=(\Sigma_{1},\dots,\Sigma_{n})$, and $\Theta:=(\Lambda,\bm{\mu},\bm{\Sigma})$.
@@
Algorithm 1 Two-step estimation via random projections and ECF
@@
2:Step 1: For each $u_{r}$, estimate $(\widehat{\upsilon}_{u_{r},(1)},\widehat{\sigma}^{2}_{u_{r},(1)},\widehat{\Lambda}_{u_{r},(1)})$ by minimizing $(Z_{u}-F_{u})^{\top}W(Z_{u}-F_{u})$; set $\widehat{\Lambda}_{(1)}=\frac{1}{k}\sum_{r}\widehat{\Lambda}_{u_{r},(1)}$
[…]
4:Reconstruction: From $\{\widehat{\upsilon}_{u_{r},(2)},\widehat{\sigma}^{2}_{u_{r},(2)}\}$ solve equations
@@
for $j=1,\dots,n$, with $S^{+,d}$ the cone of $d\times d$ positive semidefinite matrices.
@@
(vi) Lastly, we note that the algorithm can also be applied to mixtures of multivariate $t$-distributions.
@@ 3.3 A consistency result
Therefore our strong consistency result is universal, i.e., it holds without further assumptions for Gaussian or $t$-mixtures. Moreover, we only need $(1/2)(2k-1)(d^{2}+d-2)$ random directions, $k$ being the number of elements in the mixture and $d$ the dimension of the space, i.e., linear in the number of components and quadratic in the dimension. No restriction is placed on the sample size.
@@
(i) Given two $n$-dimensional GMMs of $k$ Gaussians, $F=\sum_{i}w_{i}N(\mu_{i},\Sigma_{i})$ and $\widehat{F}=\sum_{i}\widehat{w}_{i}N(\widehat{\mu}_{i},\widehat{\Sigma}_{i})$, we call $\widehat{F}$ an $\epsilon$-close estimate for $F$ if there is a permutation function $\pi:[k]\to[k]$ such that, for all $i\in[k]$,
@@
Given an $n$-dimensional mixture of $k$ Gaussians $F$ that is $\epsilon$-statistically learnable, there is an algorithm that, with probability at least $1-\delta$, outputs an $\epsilon$-close estimate $\widehat{F}$, and the running-time and data requirements of the algorithm (for any fixed $k$) are polynomial in $n$, $1/\epsilon$ and $1/\delta$.
@@ 3.4 Simulations
Using the above algorithm, we estimate the parameters of a mixture of two bivariate t-distributions $\textrm{t}\left(\eta,\mu,\Sigma\right)$ with $\eta$, $\mu$ and $\Sigma$ the degrees of freedom, location and scale matrix parameters respectively. Consider
$F$ $=\lambda_{1}\,t(\eta,\mu_{1},\Sigma_{1})+(1-\lambda_{1})\,t(\eta,\mu_{2},\Sigma_{2}),$
$\mu_{1}$ $=(1,0),\quad\mu_{2}=(5,0),$
@@
From $1000$ iid simulations of $F$, see Figure 1, the objective is to estimate the parameters $\mu_{1}$, $\mu_{2}$, $\Sigma_{1}$ and $\Sigma_{2}$. We implement the proposed algorithm by projecting the data in $k=50$ directions, chosen at random. This procedure is replicated $100$ times.
Fig. 1: Depiction of $1000$ simulations of the mixture $F$.
@@
Table 1 shows the average estimates of the parameters in the replicates with their respective deviations. Here, we can clearly observe the adequate fitting of our estimators (at least within this brief example).
Table 1: Averages values of the estimated parameters over $100$ replicates of the algorithm. Marginal standard deviations (sd) are shown in parentheses.
Algorithm Component $\mu_{1}$ $\mu_{2}$ $\sigma_{11}$ $\sigma_{22}$ $\sigma_{21}$
Average estimated values
Values 1 1 0 1 1/2 0
2 5 0 1/2 1 0
RP 1 0.99 -0.01 0.99 0.49 0
sd (0.10) (0.10) (0.14) (0.13) (0.07)
2 4.99 0.014 0.56 0.73 -0.01
sd (0.13) (0.12) (0.05) (0.19) (0.06)
EM-st 1 1.22 0.07 0.89 0.62 0.01
sd (0.15) (0.16) (0.14) (0.13) (0.09)
2 4.56 -0.10 0.51 1.15 -0.01
sd (0.10) (0.25) (0.07) (0.15) (0.07)
In each replicate, we can calculate the $L^{2}$-distance (error) between the estimated mean and mean vectors and their respective true values, see Figure 2. The estimation of the mixture parameter is similar, giving on average $\widehat{\lambda_{1}}_{EM-st}=0.312$ and $\widehat{\lambda_{1}}_{RP}=0.293$ for EM-st and RP respectively. The code shows the convergence of the EM algorithm and similar execution times in both algorithms.
Fig. 2: Comparison between the EM-st and RP methods using the $L^{2}$-distance between the parameters.
@@ 3.4.2 Example 2
We repeat the previous example, but now simulating the data for the mixture of two Gaussian distributions, contaminated with a noise from a bivariate Cauchy distribution centered at $\mu_{3}$ (which we denote by $\mathcal{C}a\left(\mu_{3}\right)$). Thus
$F$ $=\lambda_{1}\,N(\mu_{1},\Sigma_{1})+(1-\lambda_{1}-\gamma)\,N(\mu_{2},\Sigma_{2})+\gamma\,\mathcal{C}a(\mu_{3}),$
$\mu_{1}$ $=(1,0),\quad\mu_{2}=(5,0),\quad\mu_{3}=(3,0),$
$\Sigma_{1}$ $=\begin{pmatrix}4&0\\ 0&2\end{pmatrix},\quad\Sigma_{2}=\begin{pmatrix}2&0\\ 0&4\end{pmatrix}$,
[…]
Here $\gamma$ indicates the proportion of outliers. In this case we re-estimate the original mixture parameters, and the $L^{2}$-errors are plotted in Figure 3.
@@
Figure 3 shows that, in this example, our method presents in general smaller errors in the estimation of the parameters than RobEM.
Fig. 3: Comparison between the Robust EM and RP Boxplots when outliers are added, taking $\mu_{3}=(3,0)$.

### v4

@@ Abstract
We address two important statistical problems: that of estimating for mixtures of multivariate normal distributions and mixtures of $t$-distributions based on univariate projections, and that of measuring the agreement between two different random partitions. The results are based on an earlier work of the authors, where it was shown that mixtures of multivariate Gaussian or $t$-distributions can be distinguished by projecting them onto a certain predetermined finite set of lines, the number of lines depending only on the total number of distributions involved and on the ambient dimension. We also compare our proposal with robust versions of the expectation-maximization method EM. In each case, we present algorithms for effecting the task, and compare them with existing methods by carrying out some simulations.
@@ 1.1.1 RP via the Johnson–Lindenstrauss lemma
Lemma (Johnson–Lindenstrauss lemma, [46]). Let $\aleph_{n}$ be a set of $n$ points in $\mathbb{R}^{d}$, let $0<\epsilon<1$ and let $q>(8\log n)/\epsilon^{2}$. Then there exists a linear map $f:\mathbb{R}^{d}\to\mathbb{R}^{q}$ such that, for all $u,v\in\aleph_{n}$,
@@
Let $\aleph_{n}$ and $\epsilon$ be as in Theorem 1, and let $q>(4\log n)/(\epsilon^{2}/2-\epsilon^{3}/3)$. Define $f(u):=\sqrt{(d/q)}PUu$, where $U$ is a $d\times d$ unitary matrix whose columns are iid (independent and identically distributed) random vectors uniformly distributed on the unit sphere in $\mathbb{R}^{d}$, and $P:\mathbb{R}^{d}\to\mathbb{R}^{q}$ denotes the projection onto the first $q$ coordinates. Then
@@ 1.1.3 On the pros and cons of RP-methods
Although the manuscript addresses (i) parameter estimation for multivariate mixtures and (ii) agreement between random partitions, both contributions are governed by the same principle: within the Gaussian/$t$-mixture classes considered here, the underlying distribution is determined by a finite collection of one-dimensional projections along a strong sm-uniqueness set (Section 2). Section 3 exploits this fact constructively for estimation, whereas Section 4 exploits it comparatively to quantify agreement by aggregating projected discrepancies.
Notation.
Throughout the manuscript, $m$ denotes the number of mixture components, $d$ the ambient dimension, $N$ the sample size (i.i.d. observations), and $k$ the number of projection directions used by the random-projection procedures. In particular, we use $N$ for the sample size to avoid notational conflicts with other indices.
@@ 1.2 Roadmap of the rest of the article
Our arguments rely on the notion of a strong symmetric-matrix uniqueness set (strong sm-uniqueness set), which provides a finite Cramér–Wold system for Gaussian and $t$-mixtures. We recall the definition and its construction in Section 2; see [35, 34].
@@ 2.3 Strong sm-uniqueness sets
Theorems 4 and 5 beg the question as to whether there exist strong sm-uniqueness sets of arbitrarily large cardinality. An affirmative answer is given in the following result based on [34, Theorem 4.3], which also provides a realistic method for generating them.
We next recall the notion of a (strong) sm-uniqueness set, which is needed for the theorem below.
Definition 1 (sm-uniqueness and strong sm-uniqueness sets [35, 34]).
Let $S\subset\mathbb{R}^{d}$. We say that $S$ is an sm-uniqueness set if the only real symmetric $d\times d$ matrix $A$ satisfying $x^{\top}Ax=0$ for all $x\in S$ is $A=0$. If $|S|\geq d(d+1)/2$, we say that $S$ is strong if every subset of $S$ with cardinality $d(d+1)/2$ is itself an sm-uniqueness set.
@@ 3.2 Description of the algorithm
$\sum_{j=1}^{m}{\lambda_{j}}\frac{1}{(2\pi\det(\Sigma_{j}))^{1/2}}\exp\Bigl(-\tfrac{1}{2}(x-\mu_{j})^{T}\Sigma_{j}^{-1}(x-\mu_{j})\Bigr),\quad(x\in\mathbb{R}^{d}),$
where $\lambda_{1},\dots,\lambda_{m}\in[0,1]$, $\sum_{j=1}^{m}\lambda_{j}=1$, $\mu_{1},\dots,\mu_{m}\in\mathbb{R}^{d}$, and $\Sigma_{1},\dots,\Sigma_{m}$ are $d\times d$ positive-definite matrices. We denote $\Lambda:=(\lambda_{1},\dots,\lambda_{m})$, $\bm{\mu}:=(\mu_{1},\dots,\mu_{m})$, $\bm{\Sigma}:=(\Sigma_{1},\dots,\Sigma_{m})$, and $\Theta:=(\Lambda,\bm{\mu},\bm{\Sigma})$.
@@
Algorithm 1 Two-step estimation via random projections and ECF with alignment
@@
2:Step 1: For each $u_{r}$, estimate $(\widehat{\upsilon}_{u_{r},(1)},\widehat{\sigma}^{2}_{u_{r},(1)},\widehat{\Lambda}_{u_{r},(1)})$ by minimizing $(Z_{u}-F_{u})^{\top}W(Z_{u}-F_{u})$
3:Alignment: Choose a pivot direction $u_{r^{\ast}}$ (e.g., with largest separation between $\widehat{\upsilon}_{u_{r},(1),j}$), order its components by $\widehat{\upsilon}_{u_{r^{\ast}},(1),j}$, and treat the ordered triples $(\widehat{\Lambda}_{u_{r^{\ast}},(1),j},\widehat{\upsilon}_{u_{r^{\ast}},(1),j},\widehat{\sigma}^{2}_{u_{r^{\ast}},(1),j})$ as prototypes. For each $u_{r}$, solve an assignment problem (Hungarian/global arg min) that matches the triples $(\widehat{\Lambda}_{u_{r},(1),j},\widehat{\upsilon}_{u_{r},(1),j},\widehat{\sigma}^{2}_{u_{r},(1),j})$ to the pivot prototypes, obtaining aligned parameters $(\widetilde{\upsilon}_{u_{r},(1)},\widetilde{\sigma}^{2}_{u_{r},(1)},\widetilde{\Lambda}_{u_{r},(1)})$.
4:Set $\widehat{\Lambda}_{(1)}=\frac{1}{k}\sum_{r=1}^{k}\widetilde{\Lambda}_{u_{r},(1)}$
5:Reconstruction: From $\{\widehat{\upsilon}_{u_{r},(2)},\widehat{\sigma}^{2}_{u_{r},(2)}\}$ solve, for $j=1,\dots,n$,
@@
with $S^{+,d}$ the cone of $d\times d$ positive semidefinite matrices.
@@
In Algorithm 1, the parameters are initialized separately in each random one-dimensional projection and the main driver of the initialization is a $k$-means split. Specifically, for a randomly sampled direction $u$ (normalized to unit length), the data are projected as $Y_{i}=u^{\top}X_{i}$ and we run $k$-means with two clusters on $\{Y_{i}\}_{i=1}^{N}$. The resulting two clusters provide a data-driven starting point: the initial projected locations are set to the within-cluster sample means, the projected scales to the within-cluster sample variances (truncated below by a small constant for numerical stability), and the initial mixing weight is taken as the empirical cluster proportion. To avoid poor starts, we only keep projections where both clusters contain at least a minimum number of observations and the two $k$-means centroids are sufficiently separated relative to $\mathrm{sd}(Y)$. After ordering components by increasing projected mean, the $k$-means-based initialization is optionally refined by a single soft-EM update for the mixing weight, and then passed to the GEL estimation step in that projection.
The reconstruction of each covariance matrix $\Sigma_{j}$ from the projected variances is obtained by minimizing the quadratic expression
$f(\Sigma)=\sum_{r=1}^{k}\bigl(u_{r}^{\top}\Sigma u_{r}-s_{r,j}\bigr)^{2},\qquad\Sigma\in\mathbb{S}^{d}.$
This objective is always convex in $\Sigma$, but uniqueness of the solution requires strict convexity, which in turn depends on the geometry of the directions $\{u_{r}\}$. The following proposition makes this condition explicit.
Proposition 1 (Uniqueness of covariance reconstruction).
Let $u_{1},\dots,u_{k}\in\mathbb{S}^{d-1}$ and consider
$f(\Sigma)=\sum_{r=1}^{k}\bigl(u_{r}^{\top}\Sigma u_{r}-s_{r}\bigr)^{2},\qquad\Sigma\in\mathbb{S}^{d}.$
If $\{u_{r}u_{r}^{\top}:r=1,\dots,k\}$ spans $\mathbb{S}^{d}$, equivalently if $\mathrm{rank}(A)=d(d+1)/2$ where the $r$-th row of $A$ is $\mathrm{vecsym}(u_{r}u_{r}^{\top})$, then $f$ is strictly convex on $\mathbb{S}^{d}$ and admits a unique minimizer. In particular, the reconstruction equations (4)(5) have a unique solution for each covariance matrix $\Sigma_{j}$.
Sketch of proof.
Identifying $\mathbb{S}^{d}$ with $\mathbb{R}^{m}$, $m=d(d+1)/2$, via $\theta=\mathrm{vecsym}(\Sigma)$, we can write
$f(\Sigma)=\bigl\|A\,\theta-s\bigr\|_{2}^{2},$
with $A\in\mathbb{R}^{k\times m}$ having $r$-th row $\mathrm{vecsym}(u_{r}u_{r}^{\top})$. The spanning condition $\{u_{r}u_{r}^{\top}\}$ spans $\mathbb{S}^{d}$ is equivalent to $\mathrm{rank}(A)=m$, so $A^{\top}A$ is positive definite. Hence the Hessian of $f$ with respect to $\theta$ is $2A^{\top}A\succ 0$, and $f$ is strictly convex, with a unique minimizer $\theta^{\star}$, hence a unique $\Sigma^{\star}$. ∎
@@
(vi) We note that the algorithm can also be applied to mixtures of multivariate $t$-distributions.
(vii) It is important to note that Step 1 can be implemented with any other consistent univariate-mixture estimators at each direction $u_{r}$, for instance the univariate EM algorithm, instead of the ECF-based criterion.
@@ 3.3 A consistency result
Therefore our strong consistency result is universal, i.e., it holds without further assumptions for Gaussian or $t$-mixtures. Moreover, the theorem-level requirement in [34] shows that it suffices to take
$k\;\geq\;\frac{1}{2}(2m-1)(d^{2}+d-2)+1$
projection directions, where $m$ denotes the number of mixture components and $d$ the ambient dimension. Hence the required number of directions grows linearly in $m$ and quadratically in $d$.
No restriction is placed on the sample size.
@@
(i) Given two $d$-dimensional GMMs of $m$ Gaussians, $F=\sum_{i}w_{i}N(\mu_{i},\Sigma_{i})$ and $\widehat{F}=\sum_{i}\widehat{w}_{i}N(\widehat{\mu}_{i},\widehat{\Sigma}_{i})$, we call $\widehat{F}$ an $\epsilon$-close estimate for $F$ if there is a permutation function $\pi:[k]\to[k]$ such that, for all $i\in[k]$,
@@
Given an $d$-dimensional mixture of $m$ Gaussians $F$ that is $\epsilon$-statistically learnable, there is an algorithm that, with probability at least $1-\delta$, outputs an $\epsilon$-close estimate $\widehat{F}$, and the running-time and data requirements of the algorithm (for any fixed $m$) are polynomial in $n$, $1/\epsilon$ and $1/\delta$.
@@ 3.4 Simulations
Using the above algorithm, we estimate the parameters of a mixture of two bivariate $t$-distributions $t(\nu,\mu,\Sigma)$, where $\nu$, $\mu$, and $\Sigma$ denote the degrees of freedom, location, and scale matrix parameters, respectively. Consider
$F$ $=\lambda_{1}\,t(\nu,\mu_{1},\Sigma_{1})+(1-\lambda_{1})\,t(\nu,\mu_{2},\Sigma_{2}),$
$\mu_{1}$ $=(0,0),\quad\mu_{2}=(\eta,0),$
@@
Based on $200$ i.i.d. draws from $F$ (see Figure 1), our goal is to estimate $\mu_{1}$, $\mu_{2}$, $\Sigma_{1}$, and $\Sigma_{2}$. We implement the proposed algorithm by projecting the data onto $k=50$ randomly selected directions. This experiment is replicated $100$ times. In addition, we consider four separability scenarios by letting $\eta\in\{1/2,\,1,\,3/2,\,2\}$, which progressively increases the separation between the component locations.
Fig. 1: Bivariate $t$-mixture samples for increasing separation $\eta$ (four panels: $\eta\in\{1/2,1,3/2,2\}$).
@@
Table 1 reports the average parameter estimates across the replicates, together with their corresponding standard deviations. In addition, we report the (replicate-averaged) confusion matrices (see Table 2) computed over the $100$ Monte-Carlo runs—comparing the true component labels with the posterior (MAP) allocations obtained after parameter estimation, for both the EM-st and the RP procedures. These confusion matrices are provided in Appendix B.
In each replicate, we compute estimation errors by comparing the estimated parameters with their true values. Specifically, we construct boxplots of (i) the $L^{2}$-error for the mixing weight $\lambda_{1}$, (ii) the $L^{2}$-error for the mean vectors, and (iii) the Frobenius-distance for the variance–covariance matrices. These boxplots are reported for each separability scenario $\eta\in\{1/2,1,3/2,2\}$; see Figure 3 for EM-st and RP, respectively. Overall, both algorithms exhibit comparable performance across scenarios. In Figure 2, the boxplots show that the estimation errors for the mixing coefficients $\lambda$ are similar for both methods.
Fig. 2: Mixing-weight error for $\lambda_{1}$: EM-st vs RP across separability scenarios $\eta\in\{1/2,1,3/2,2\}$ (boxplots over 100 replicates).
Fig. 3: Comparison between the EM-st (EM) and RP methods via boxplots of estimation errors across the four separability scenarios $\eta\in\{1/2,1,3/2,2\}$. Errors are measured using the $L^{2}$-distance for the mean vectors, and the Frobenius distance for the variance–covariance matrices.
@@ 3.4.2 Example 2
We repeat the previous simulation setup, but now the observations are generated from a two-component bivariate Student $t$-mixture with $\nu=4$ degrees of freedom, contaminated by a fraction $\gamma$ of uniform noise on an axis-aligned square. Specifically, for each replicate we generate $N=500$ observations from
$F=(1-\gamma)\Big\{\lambda_{1}\,t_{\nu}(\mu_{1},\Sigma_{1})+\lambda_{2}\,t_{\nu}(\mu_{2},\Sigma_{2})\Big\}+\gamma\,\mathcal{U}\!\big([0,4]\times[0,4]\big),$
where $\lambda_{1}=0.3$ and $\lambda_{2}=0.7$, and
$\mu_{1}=(0,0),\qquad\mu_{2}=(2,0),$
$S_{1}=\begin{pmatrix}1&0\\[1.00006pt] 0&1/2\end{pmatrix},\qquad S_{2}=\begin{pmatrix}1/2&0\\[1.00006pt] 0&1\end{pmatrix}$.
The contamination distribution is uniform on the square centered at $(2,2)$ with side length $4$, i.e., $\mathcal{U}([0,4]\times[0,4])$. The parameter $\gamma$ controls the proportion of outliers, and we consider $\gamma\in\{0.05,\,0.10,\,0.15\}$. Figure 4 displays one simulated dataset for each noise scenario considered in our study. Each panel corresponds to a different contamination level $\gamma\in\{0.05,0.10,0.15\}$ and illustrates a single realization from the bivariate Student $t$-mixture with uniform outliers; hence, the figure provides a visual summary of how the amount of uniform noise increases across scenarios.
Fig. 4: Simulated bivariate Student $t$-mixture with uniform contamination. Each panel shows a sample of size $N=500$ from a two-component $t_{\eta}$ mixture with $\eta=4$, mixing weights $(\lambda_{1},\lambda_{2})=(0.3,0.7)$, locations $\mu_{1}=(0,0)$ and $\mu_{2}=(2,0)$, and scatter matrices $\Sigma_{1}=\mathrm{diag}(1,1/2)$ and $\Sigma_{2}=\mathrm{diag}(1/2,1)$. A proportion $\gamma\in\{0.05,0.10,0.15\}$ of observations is replaced by outliers drawn uniformly from $0,4]\times[0,4]$.
In this case we re-estimate the original mixture parameters, and the $L^{2}$-errors are plotted in Figure 5.
@@
In the robust contamination setting, the RP approach exhibits an overall better performance than the robust EM alternative. This improvement is particularly clear for the mixing proportion parameter $\lambda$, whose estimation is more stable under RP across the noise scenarios, as well as for the scatter parameters: the RP reconstructions of the variance–covariance (scatter) matrices show smaller discrepancies with respect to the true matrices, indicating a higher robustness of RP to the uniform outliers when recovering second-order structure, see Figure 5 and Figure 6.
Fig. 5: Comparison of estimation errors for the mixing proportion $\lambda_{1}$ under uniform contamination. Each panel corresponds to a different outlier proportion $\gamma\in\{0.05,0.10,0.15\}$ (left to right), and boxplots summarize the errors over $100$ Monte Carlo replicates for the RobEM and RP methods.
Fig. 6: Comparison of estimation errors for the remaining parameters under uniform contamination. Each panel corresponds to a different outlier proportion $\gamma\in\{0.05,0.10,0.15\}$ (left to right), and boxplots summarize the errors over $100$ Monte Carlo replicates for the RobEM and RP methods. Errors are measured using the $L^{2}$-distance for the mean vectors and the Frobenius distance for the variance–covariance matrices.
Table 3 reports the average parameter estimates across the replicates, together with their corresponding standard deviations, under the contamination setting with uniform outliers. In addition, we report the (replicate-averaged) confusion matrices (see Table 4) computed over the $100$ Monte-Carlo runs—comparing the true component labels with the posterior (MAP) allocations obtained after parameter estimation, for both the robust EM and the RP procedures, see Appendix B.